# Digital Network Twins for Next-generation Wireless: Creation, Optimization, and Challenges

### v1

@@ IV Case Studies
This section showcases the creation and optimization of DNTs in two case studies: wireless traffic forecasting and reliable edge caching, with respect to the proposed methods in Sec. II-A and II-B.
@@ IV-A Joint Vertical-Horizontal Twinning for Traffic Prediction
Specifically, we employ the real-world network traffic dataset [15] of Milan City, which is composed of urban telecommunications behaviors intricately divided into 10,000 grid cells, with each cell equipped with a dedicated BS that captures a wealth of time-series data records, including detailed information on short message service (SMS), call activities, and internet usage patterns. We adopt mean squared error (MSE), mean absolute error (MAE), and normalized root mean square error (NRMSE) as key performance metrics to measure the distinctions between ground-truth and generated data from the mapped DNTs. The ground truth data are time-series real-world traffic data collected from the physical cellular network. Another key performance metric is the cost of the mapping process, which includes both training and communication time. In traditional DNT, all mapping is performed on a centralized server, requiring raw data to be uploaded for one-time modeling, which significantly increases deployment costs compared to our joint V-H twin method with continuous adjustments. H-twin incurs higher costs compared to V-twin, as DNTs demand continuous updates over extended periods rather than just initial construction. For comparison, we skip the BS clustering stage to create two twinning baselines, namely, V-twin w/o. cluster. and H-twin w/o. cluster. As shown in Fig. 4 with an example of SMS traffic data, all four twinning methods provide comparable prediction error rates, with the H-twin achieving slightly lower prediction errors. However, H-twin needs to aggregate local twin models from the BSs and dispatch them back, thereby slightly increasing the communication cost. Notably, clustering of BSs can significantly reduce the sum of communication and training costs by approximately 43.8% for H-twin. Despite the overhead introduced by H-twin, its asynchronous twinning process can effectively handle real-time heterogeneous wireless data, resulting in a reliable and accurate DNT as the time-series data evolves.
@@ IV-B DNT-Assisted Reliable Edge Caching
Here, we consider a cellular wireless network consisting of five BSs. Each BS is equipped with a local cache unit with a capability of 150 cache slots. A BS can only access its local cache unit and is forbidden to retrieve any cache units from its neighboring BSs. Each BS provides service for eight clients. The service can be overlapped, e.g., one client can be served by up to two BSs. We then create a DNT via vertical mapping from historical caching data, including information about user requests, contents, frequency, etc. The DNT serves as a data generator, providing ample datasets for training and predictive evaluation of different cache replacement policies. It produces one-step forecasts for the subsequently requested content. Moreover, various sets of historical caching data can be employed to feed the DNT, generating datasets that cover both common and rare wireless scenarios. We then implement four edge caching systems with or without the DNT. 1) RL, a basic reinforcement learning model that uses data collected from physical cellular networks for training and prediction solely; 2) RL w/ DNT, the model trained with data from both BSs and the virtual DNT; 3) Reliable RL, the model augmented by integrating safety state, action and reward modules to monitor and reduce the load imbalance among BSs; 4) Reliable RL w/ DNT, the reliable RL model trained with diverse data from both BSs and the virtual DNT. Figs. 5 and 6 evaluate the performance of four systems in terms of caching hit rate, number of risk interventions (introduced by safe modules from reliable RL), and BS load. The inclusion of DNT with safety modules achieves a higher hit rate with balanced BS loads, hence enhancing the reliability and efficacy in managing cached content across densely deployed small-cell BSs. The use of DNT also reduces the intervention from the elaborated safety constraints, suggesting that the augmented data from DNTs can improve the robustness of RL models with fewer interventions.
V Conclusion
@@
This article presents a broad overview of DNTs and unveils the vast potential of their integration for nextG wireless networks. We start by introducing the basic concepts and fundamental design of DNTs, with three highlighted applications including what-if analysis and network planning. We then describe the critical research challenges and opportunities of DNTs, in the aspects of creation, adaptation, optimization, and security. Lastly, we showcase the use of DNTs in two wireless scenarios: traffic forecasting and edge caching. We present a variety of research opportunities in addressing the challenges of DNTs, with the goal of fully harnessing their potential in future communication networks.

### v2

@@ IV Case Studies
This section showcases the creation and optimization of DNTs in three network case studies: wireless traffic forecasting, reliable edge caching, and secure DNT evolution, with respect to the proposed methods in Sec. III. In these case studies, we explicitly address the four primary challenges outlined for constructing and deploying DNTs: scalability, adaptability, robustness, and security. First, the wireless-traffic forecasting scenario leverages hierarchical twin architectures (V-twin and H-twin) combined with both synchronous and asynchronous FL and strategic BS clustering, effectively addressing scalability and adaptability by significantly reducing modeling overhead and rapidly updating models in response to dynamic traffic conditions. Second, the reliable edge-caching scenario employs DNT-based environments that enable safe reinforcement learning by simulating rare and hazardous scenarios offline, thereby directly tackling the robustness challenge by minimizing the operational risk during policy training. Lastly, the case study on traffic poisoning attacks explicitly confronts the security challenge by introducing a Twin Inconsistency Defense (TID) scheme, which statistically trims anomalous model parameters to maintain attribute aggregation integrity under sophisticated attacks. Collectively, these case studies demonstrate practical strategies to address each challenge, highlighting the effectiveness and versatility of DNTs.
@@ IV-A Joint Vertical-Horizontal Twinning for Traffic Prediction
Specifically, we employ the real-world network traffic dataset [16] of Milan City, which is composed of urban telecommunications behaviors intricately divided into 10,000 grid cells, with each cell equipped with a dedicated BS that captures a wealth of time-series data records, including detailed information on short message service (SMS), call activities, and internet usage patterns. We adopt mean squared error (MSE), mean absolute error (MAE), and normalized root mean square error (NRMSE) as key performance metrics to measure the distinctions between ground-truth and generated data from the mapped DNTs. The ground truth data are time-series real-world traffic data collected from the physical cellular network. Another key performance metric is the cost of the mapping process, which includes both training and communication time. In traditional DNT, all mapping is performed on a centralized server, requiring raw data to be uploaded for one-time modeling, which significantly increases deployment costs compared to our joint V-H twin method with continuous adjustments. H-twin incurs higher costs compared to V-twin, as DNTs demand continuous updates over extended periods rather than just initial construction. For comparison, we skip the BS clustering stage to create two twinning baselines, namely, V-twin w/o. cluster. and H-twin w/o. cluster. As shown in Fig. 3 with an example of SMS traffic data, all four twinning methods provide comparable prediction error rates, with the H-twin achieving slightly lower prediction errors. However, H-twin needs to aggregate local twin models from the BSs and dispatch them back, thereby slightly increasing the communication cost.
While VH-Twin offers only modest gains during the initial, one-off V-twin phase, its advantages become evident in the subsequent H-twin cycle. The horizontal stage delivers higher prediction accuracy than a Traditional DT approach because asynchronous parameter updates assimilate the continuous, heterogeneous data stream, keeping the twin both stable and precise. It also cuts communication and computation costs by about 43.8% relative to the baseline, single-level scheme. Taken together—slightly longer first-time initialization offset by far cheaper ongoing synchronization—the combined vertical-plus-horizontal workflow provides a more resource-efficient solution for building and maintaining DNTs.
@@ IV-B DNT-Assisted Reliable Edge Caching
Here, we consider a cellular wireless network consisting of five BSs. Each BS is equipped with a local cache unit with a capability of 150 cache slots. A BS can only access its local cache unit and is forbidden to retrieve any cache units from its neighboring BSs. Each BS provides service for eight clients. The service can be overlapped, e.g., one client can be served by up to two BSs. We then create a DNT via vertical mapping from historical caching data, including information about user requests, contents, frequency, etc. The DNT serves as a data generator, providing ample datasets for training and predictive evaluation of different cache replacement policies. It produces one-step forecasts for the subsequently requested content. Moreover, various sets of historical caching data can be employed to feed the DNT, generating datasets that cover both common and rare wireless scenarios. We then implement four edge caching systems with or without the DNT. 1) RL, a basic reinforcement learning model that uses data collected from physical cellular networks for training and prediction solely; 2) RL w/ DNT, the model trained with data from both BSs and the virtual DNT; 3) Reliable RL, the model augmented by integrating safety state, action and reward modules to monitor and reduce the load imbalance among BSs; 4) Reliable RL w/ DNT, the reliable RL model trained with diverse data from both BSs and the virtual DNT. Figs. 4 and 5 evaluate the performance of four systems in terms of caching hit rate, number of risk interventions (introduced by safe modules from reliable RL), and BS load. The inclusion of DNT with safety modules achieves a higher hit rate with balanced BS loads, hence enhancing the reliability and efficacy in managing cached content across densely deployed small-cell BSs.
Conventional data-driven optimizers often overlook decision-making reliability, leaving BSs prone to overload, uneven traffic distribution, and poor quality of experience. D-REC remedies this gap by introducing control modules into a constrained Markov decision process. These modules continually reshape the action, reward, and state definitions to respect explicit reliability targets, thereby curbing the risk of large-scale failures. Here, DNTs act as RL co-processors and safety monitors, mining diverse traffic traces to forecast the impact of cache-replacement moves in dense small-cell deployments. The use of DNT also reduces the intervention from the elaborated safety constraints, suggesting that the augmented data from DNTs can improve the robustness of RL models with fewer interventions.
IV-C Twin Poisoning Attacks to DNTs
Accurate DNT application is increasingly critical as device counts and data-driven applications surge, yet distributed learning frameworks that build our DNTs are vulnerable to twin poisoning attacks. The attackers have the ability to subtly manipulate local twin model updates and further degrade network control and management. Existing defenses focus mainly on classification tasks, leaving regression-oriented DNT workloads, where adversaries manipulate continuous time-series inputs. We introduce an attack on distributed DNT-assisted wireless networks, namely Traffic Poisoning Injection (TPI), that forgoes hard-to-breach BSs by deploying low-cost fake DNTs. It uses only each attacker’s initial model and the current global DNT to construct fake but misleading local twin models to bypass the detection from the global DNT. To mitigate this threat, we present a Twin Inconsistency Defense (TID) strategy, which statistically trims dimension-wise outlier parameters and then re-weights the benign DNT models during the attribute aggregation process.
In this part, we also consider the same real-world network traffic prediction setting in Sec. IV-A. We benchmark our TPI attack against several state-of-the-art traffic poisoning attacks and defenses to underscore its effectiveness. Additionally, we employ these baseline attacks to demonstrate the efficacy of our TID defense strategy. TABLE II illustrates that TPI incapacitates every aggregation rule tested. For Mean and FLTrust, both MAE and MSE score over the 100-point cap in the construction (V-twin) and maintenance (H-twin) phases, meaning the predicted results deviate greatly from ground truth. Even Median, which is unaffected by the MPAF baseline attack, collapses under TPI, increasing from an MAE of 0.281 to 100.00 during V-twin and from 0.296 to 100.00 during H-twin. By contrast, MPAF degrades only Mean and FLTrust, leaving Median and TID essentially unchanged. Across all scenarios, TID is the only method that remains accurate. Both MAE and MSE remain at 0.281/0.107 during V-twin and 0.27/0.10 during H-twin under MPAF attacks, while TPI raises errors to 72.5/27.5, lower than the hard-failure level of competing rules. These results establish TPI as a potent, rule-agnostic poisoning strategy and TID as the most robust defense throughout the DNT lifecycle.
TABLE II: Performance on the Milan‐Internet Dataset for Both V-twin and H-twin.
Defense Metric V-twin H-twin
NO MPAF FTI NO MPAF TPI
Mean MAE 0.281 100.0 100.0 0.266 100.0 100.0
MSE 0.106 100.0 100.0 0.101 100.0 100.0
Median MAE 0.281 0.281 100.0 0.296 0.296 100.0
MSE 0.106 0.106 100.0 0.101 0.101 100.0
FLTrust MAE 0.312 100.0 100.0 0.297 100.0 100.0
MSE 0.114 100.0 100.0 0.109 100.0 100.0
TID MAE 0.281 0.281 72.453 0.266 0.266 72.458
MSE 0.106 0.107 27.548 0.101 0.102 27.543
IV-D Lessons Learned & Discussion
The first case study explores wireless traffic forecasting in a city-scale cellular network covering approximately 10,000 grid-based cells within Milan. Here, DNT functions primarily as a predictive analyzer, integrating both historical traffic patterns (V-twin) and incremental, localized updates via asynchronous federated learning (H-twin). Such scenarios, characterized by abundant spatial-temporal data and relatively low direct-testing risk, highlight DNTs’ strengths in efficiently tracking and predicting dynamic network states, making them particularly suitable for network management and proactive resource provisioning tasks. The second case study addresses reliable edge-caching in a small-cell cluster, consisting of five base stations with overlapping coverage and stringent QoS constraints. Here, the DNT acts as a critical sandbox and safety-net environment, synthesizing rare and potentially harmful network conditions to safely guide RL-based caching policies. Scenarios such as this, marked by high experimentation risks, tight latency requirements, and the need to safely explore counterfactual and rare-event scenarios, clearly benefit from the DNT’s capability to rigorously test policies offline prior to deployment. In the third case study, we explore the vulnerability of distributed DNTs to model poisoning attacks within a real-world traffic prediction setting. This scenario emphasizes that applications involving decentralized learning with continuous and dynamic data inputs particularly benefit from robust anomaly-detection defenses embedded within the DNT framework.
V Conclusion and Future Directions
@@
This article presents a broad overview of DNTs and unveils the vast potential of their integration for nextG wireless networks. We start by introducing the basic concepts and fundamental design of DNTs, with three highlighted applications, including what-if analysis and network planning. We then describe the critical research challenges and opportunities of DNTs, in the aspects of creation, adaptation, optimization, and security. Lastly, we showcase the use of DNTs in two wireless scenarios: traffic forecasting and edge caching. We present a variety of research opportunities in addressing the challenges of DNTs, with the goal of fully harnessing their potential in future communication networks.
Looking forward, several promising research directions can enhance both the capabilities and practical deployment of DNTs. First, developing automated mechanisms for dynamic twin adjustment, where DNTs intelligently balance model fidelity and computational efficiency based on real-time network states, could further optimize predictive performance. Additionally, extending DNT architectures to seamlessly integrate heterogeneous network elements (e.g., satellite, drone-based, and terrestrial networks) will broaden their applicability, such as transferring existing DNTs to non-terrestrial network scenarios. From a security perspective, designing advanced, adaptive defenses against poisoning attacks for continuously streaming data remains critical. On the deployment front, creating standardized interfaces and integration frameworks for open RAN with xApps/rApps and core networks will facilitate broader adoption and interoperability of DNTs.